THE ROTATIONAL SPECTRUM OF H_{2}S: THE H_{2}^{33}S ISOTOPOLOGUE AND THE SUB-DOPPLER RES-OLUTION IN THE THz REGIME

GABRIELE CAZZOLI, \underline{CRISTINA PUZZARINI}, Dipartimento di Chimica “G. Ciamician”, Università di Bologna, I-40126 Bologna, Italy.

Measurements of the rotational spectrum of H_{2}^{33}S, detected in natural abundance, were performed in the 200 GHz - 1.072 THz frequency range. The Lamb-dip technique was employed to exploit sub-doppler resolution and thus to resolve the hyperfine structure of the rotational lines. The retrieved transition frequency values allowed us to revise the ^{33}S spin-rotation tensor as well as to improve the ^{33}S quadrupole-coupling and centrifugal-distortion constants. Furthermore, the rotational spectrum of the main isotopopologue was investigated in the THz frequency regime. The Lamb-dip technique was employed as well in order to obtain sub-doppler resolution in this frequency region and thus to provide transition frequency values at 1 THz with an accuracy of 1 kHz.